# The Hilbert-space structure of free fermions in disguise

### v2

@@ 1 Introduction
It has been almost a hundred years since the Jordan-Wigner transformation (JW) was introduced in theoretical physics [1]. Today, this transformation remains an invaluable tool, allowing us to map one-dimensional spin systems into fermionic ones, with several generalizations being proposed over the years [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Importantly, the JW transformation and its generalizations make it easy to identify and construct exactly-solvable spin chains, which are the ones that are mapped to non-interacting fermionic Hamiltonians [18].
@@
Given the long history of the JW transformation, it is perhaps surprising that only recently new types of spin chains that can be mapped to free fermions were discovered. In 2019, Paul Fendley constructed a model which is quartic in the JW fermions, and still quadratic in terms of suitably defined fermionic creation and annihilation operators [19] (see Refs. [20, 21, 22] for earlier work in this direction). Remarkably, it was proven that the spin chain found by Fendley cannot be diagonalized by any generalized JW transformation [23], making the model fundamentally different from the previously known ones. This model and later generalizations [24, 25, 26, 27, 28] are now identified with the name of free fermions in disguise (FFD).
@@
The mapping to free fermions in FFD models gives access to the full Hamiltonian spectrum. Still, the computation of spin correlation functions is generally hard, because the dictionary between states in the spin and free-fermion Hilbert spaces is highly non-trivial. Indeed, on the one hand the FFD mapping is non-linear and non-local; on the other hand, each Hamiltonian eigenspace is known to be exponentially degenerate [19], so that fermionic Fock states cannot be uniquely associated with spin states. As a consequence, the problem of computing correlation functions in FFD models remains essentially open, despite some recent progress in addressing the “inverse problem" of expressing local spin operators in terms of fermionic ones [29]. In some respect, the current situation is similar to that of interacting integrable models [30, 31]: while the Bethe ansatz gives direct access to the Hamiltonian spectrum, computing correlation functions in arbitrary eigenstates is hard, and analytic results exist only in a few special cases [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
@@
The picture is very different for JW free fermions, where a number of simplifications occur. First, the JW transformation maps local bilinear fermionic operators into local spin operators. Second, it is often easy to characterize the family of spin states that map to the fermionic Gaussian states [^1] We recall that Gaussian states are defined as those satisfying the Wick theorem [45, 46].. These facts have consequences for the complexity of the JW solvable models. For instance, the dynamics of spin correlation functions can be computed either analytically or numerically efficiently in many cases of interest [47, 48, 49], and especially in the case of quantum quenches [50, 51, 52, 53]. In addition, the eigenstate wave functions are simple compared to generic Hamiltonians. In particular, all eigenstates can be prepared efficiently by a quantum computer, as they can be obtained as the output of a quantum circuit whose depth scales only linearly in the system size [54, 55, 56, 57, 58]. It is natural to ask whether the same conclusions can be drawn for FFD models. Can the system dynamics be simulated on a classical computer? Can all eigenstates be prepared efficiently on a quantum computer? [59] These questions are non-trivial, due to the complicated structure of the dictionary between spin and fermion degrees of freedom.
@@
In this work, we take a step towards answering these questions, by characterizing the Hilbert space associated to FFD Hamiltonians. For concreteness, we will focus on the original model introduced by Paul Fendley [19], although similar constructions hold for more general FFD Hamiltonians. We show that the corresponding Hilbert space admits the exact factorization $\mathcal{H}=\mathcal{H}_{F}\otimes\mathcal{H}_{D}$, where $\mathcal{H}_{F}$ hosts the fermionic operators, while $\mathcal{H}_{D}$ accounts for the exponential degeneracy of the energy eigenspaces. By constructing a family of spin operators generating the operator algebra supported on $\mathcal{H}_{D}$, we further show that $\mathcal{H}_{D}=\mathcal{H}_{F^{\prime}}\otimes\mathcal{H}_{\widetilde{D}}$, where $\mathcal{H}_{F^{\prime}}$ hosts ancillary free fermions in disguise, while $\mathcal{H}_{\widetilde{D}}$ is generated by the common eigenstates of an extensive set of commuting Pauli strings. Our construction allows us to fully resolve the exponential degeneracy of all Hamiltonian eigenspaces and is expected to have implications for the computation of spin correlation functions, both in and out of equilibrium.
@@
This work is organized as follows. We begin in Sec. 2 by recalling the model introduced by Paul Fendley (Sec. 2.1) and its exact solution in terms of free fermions (Sec. 2.2). Our results are contained in Secs. 2.3, 4, and 5, and summarized in Sec 3. In Sec. 2.3 we show that the Hilbert space factorizes as $\mathcal{H}=\mathcal{H}_{F}\otimes\mathcal{H}_{D}$, where $\mathcal{H}_{F}$ supports the fermionic operators. The structure of $\mathcal{H}_{D}$ is further resolved in the later sections, where we show that $\mathcal{H}_{D}=\mathcal{H}_{F^{\prime}}\otimes\mathcal{H}_{\widetilde{D}}$ and characterize the two factors individually. In particular, in Sec. 4 we show that $\mathcal{H}_{\widetilde{D}}$ is a “representation-dependent” subspace, generated by the common eigenstates of an extensive set of commuting Pauli strings, while in Sec. 5 we show that the space of operators supported on $\mathcal{H}_{F^{\prime}}$ is generated by an operator algebra of ancillary free fermion in disguise. Finally, Sec. 6 contains our conclusions and an outlook for future work.
2 The Model and its Hilbert-space structure
@@
2.1 The Hamiltonian
We first introduce the FFD Hamiltonian presented in Ref. [19]. The construction starts with an abstract operator algebra defined in terms of a set of generators $\{h_{m}\}_{m=1}^{M}$. They satisfy the commutation relations
@@ 2.1 The Hamiltonian
Different representations of this algebra give rise to different models. In this work, we focus on the original representation considered in Ref. [19], which is associated with a one-dimensional chain of $M$ spins. The Hilbert space of the model is $\mathcal{H}=(\mathbb{C}^{2})^{\otimes M}$, while
@@
where $b_{m}\in\mathbb{R}$ are arbitrary real numbers parametrizing the model (note that the Hamiltonian (2) corresponds to open boundary conditions). In the next subsection, we review the exact solution of the Hamiltonian (2). We will focus on the aspects that are directly relevant to our work, and refer to Ref. [19] for more detail.
2.2 Exact solution
The diagonalization of the Hamiltonian (2) is achieved by constructing suitable fermionic creation and annihilation operators. A key object is the transfer matrix $T_{M}(u)$, an operator acting on $\mathcal{H}$ that depends on a complex parameter $u$, called the spectral parameter. The transfer matrix can be constructed as the generating function of the conserved quantities of the model [19], in a fashion which is reminiscent of Bethe-ansatz solvable models [30].
@@ 2.2 Exact solution
As shown in [19], the creation/annihilation operators associated with the energy $\epsilon_{k}$ then take the form
$\Psi_{k}\propto T_{M}(\epsilon_{k})~\chi~T_{M}(-\epsilon_{k})\,.$ (7)
@@
2.3 The Hilbert-space structure
In this section, we use the fermionic operators (7) to show that the FFD Hilbert space can be decomposed as $\mathcal{H}=\mathcal{H}_{F}\otimes\mathcal{H}_{D}$. Here, $\mathcal{H}_{F}$ is a space of dimension $2^{S}$ that the fermionic operators are supported on, while $\mathcal{H}_{D}$ accounts for the exponentially large degeneracies observed in the spectrum of $H$. This simple observation is our first result and the basis for the subsequent analysis presented in the rest of this work, cf. Sec. 3
@@ 2.3 The Hilbert-space structure
Next, we claim that the isomorphism $\varphi$ maps the operators $\Psi_{k}$ into operators supported on $\mathcal{H}_{F}$, namely $\varphi(\Psi_{k})=c_{k}\otimes\mathbb{1}$. Indeed, by definition of $\varphi$,
@@
so that all fermionic operators are supported on $\mathcal{H}_{F}$. Finally, since the Hamiltonian is written explicitly in terms of the fermionic operators, it follows that (with a slight abuse of notation) $H=H\otimes\mathbb{1}$. Therefore, $\mathcal{H}_{D}$ accounts for the exponential degeneracy of each Hamiltonian eigenstate.
3 Overview of our results
@@
  • We will show that ${\mathcal{H}}_{F^{\prime}}$ supports the action of all operators written in terms of the generators $h_{m}$ that commute with the fermion creation/annihilation operators. We will denote the algebra of such operators by $\mathcal{A}_{F^{\prime}}$ and we will refer to it as the ancillary fermionic algebra (to be distinguished by the fermionic algebra $\mathcal{A}_{F}$ supported on $\mathcal{H}_{F})$. Importantly, the definition of $\mathcal{A}_{F^{\prime}}$ is independent of the specific representation chosen, because it can be stated entirely in terms of the generators $h_{m}$ and their algebraic relations. However, we will show that $\mathcal{A}_{F^{\prime}}$ does depend on the Hamiltonian couplings $\{b_{m}\}$. Going further, we will show in Sec. 5, that $\mathcal{A}_{F^{\prime}}$ is a fermionic algebra, generated by a new family of ancillary fermion operators
@@
The dimensions of the subspaces $\mathcal{H}_{F}$, $\mathcal{H}_{F^{\prime}}$, and $\mathcal{H}_{\widetilde{D}}$ depend, respectively, on the integers $S$, $S^{\prime}$, $N_{P}$ and $N_{C}$ introduced above. In Table 1, we summarize their dependence on the system size $M$. In all cases (except for the case $M=6k+3$, where an additional zero mode has to be included to the construction, cf. the end of Sec. 5), we can check that the product of the dimension of the subspaces $\mathcal{H}_{F}$, $\mathcal{H}_{F^{\prime}}$, and $\mathcal{H}_{\widetilde{D}}$ is equal to the dimension of the full Hilbert space, ${\rm dim}(\mathcal{H})=2^{N}$.
@@
$M$  $S=\lfloor\frac{M+2}{3}\rfloor$   $N_{C}$   $N_{P}$ $S^{\prime}=\lfloor\frac{M+2}{6}\rfloor$
@@
It is important to compare our results with some constructions found in Ref. [19]. There, a family of operators $E^{(n)},O^{(n)}$ is presented, commuting with the Hamiltonian and generating an extended supersymmetry (SUSY) algebra. While this symmetry also accounts for the degeneracies of the spectrum, it has a few drawbacks. First, it is not clear what the underlying algebra is. Second, the generators $E^{(n)},O^{(n)}$ mix with the fermions: namely, they entangle the spaces $\mathcal{H}_{F}$ and $\mathcal{H}_{D}$, which we believe makes it difficult to understand the corresponding algebra  [^2] In our construction, we will also encounter operators $E^{(n)},O^{(n)}$. Those are defined much similarly to Fendley’s, but with the important difference that their relation with the fermionic operators is clearly understood..
4 Representation-dependent degeneracies: the subspace $\mathcal{H}_{\widetilde{D}}$
As pointed out in Ref. [19], the representation (1) allows for a family of operators $\{\widetilde{h}_{m}\}$ that commute simultaneously with all the $h_{m}$, and therefore generate a symmetry of the model which does not depend on the Hamiltonian couplings $\{b_{m}\}$. More generally, such symmetry extends to any type of quantum dynamics defined in terms of the operators $\{h_{m}\}$, such as the circuits considered in Refs. [27, 60]. The operators $\widetilde{h}_{m}$ can be chosen as [19]
@@ 4 Representation-dependent degeneracies: the subspace $\mathcal{H}_{\widetilde{D}}$
each pair $(\tilde{X}_{n},\tilde{Y}_{n})$ accounting for a two-fold degeneracy of the Hamiltonian spectrum. For some values of $M$ modulo $6$, these two families are completed by either one or two additional central operators, commuting with all $\tilde{X}_{m}$ and $\tilde{Y}_{m}$. Of these central operators, some belong exclusively to the algebra of the $\widetilde{h}_{m}$, while others can also be formulated as elements of the algebra of the $h_{m}$ (see paragraphs 4.5 and 4.6): for reasons that will be made clear below, the index $N_{C}$ in Table 1 counts only the former.
4.1 $M=6k$

### v4

@@ I Introduction
It has been almost a hundred years since the Jordan-Wigner transformation (JW) was introduced in theoretical physics [jordan1993]. Today, this transformation remains an invaluable tool, allowing us to map one-dimensional spin systems into fermionic ones, with several generalizations being proposed over the years [fradkin1989jordan, wigner1991, huerta1993bose, batista2001generalized, verstraete2005mapping, kitaev2006anyons, nussinov2012arbitrary, chen2018exact, backens2019jordan, tantivasadakarn2020jordan, chapman2020characterization, minami2016solvable, minami2017infinite, yanagihara2020exact, ogura2020geometric, wang2025particle]. Importantly, the JW transformation and its generalizations make it easy to identify and construct exactly-solvable spin chains, which are the ones that are mapped to non-interacting fermionic Hamiltonians [franchini2017introduction].
@@
Given the long history of the JW transformation, it is perhaps surprising that only recently new types of spin chains that can be mapped to free fermions were discovered. In 2019, Paul Fendley constructed a model which is quartic in the JW fermions, and still quadratic in terms of suitably defined fermionic creation and annihilation operators [fendley2019free] (see Refs. [fendley2007cooper, de2016integrable, feher2019curious] for earlier work in this direction). Remarkably, it was proven that the spin chain found by Fendley cannot be diagonalized by any generalized JW transformation [elman2021free], making the model fundamentally different from the previously known ones. This model and later generalizations [alcaraz2020free, alcaraz2020integrable, fendley2024free, fukai2025quantum, fukai2025free, chapman2023unified] are now identified with the name of free fermions in disguise (FFD).
@@
The mapping to free fermions in FFD models gives access to the full Hamiltonian spectrum. Still, the computation of spin correlation functions is generally hard, because the dictionary between states in the spin and free-fermion Hilbert spaces is highly non-trivial. Indeed, on the one hand the FFD mapping is non-linear and non-local; on the other hand, each Hamiltonian eigenspace is known to be exponentially degenerate [fendley2019free], so that fermionic Fock states cannot be uniquely associated with spin states. As a consequence, the problem of computing correlation functions in FFD models remains essentially open, despite some recent progress in addressing the “inverse problem” of expressing local spin operators in terms of fermionic ones [vona2014exact]. In some respect, the current situation is similar to that of interacting integrable models [korepin1997quantum, essler2005one]: while the Bethe ansatz gives direct access to the Hamiltonian spectrum, computing correlation functions in arbitrary eigenstates is hard, and analytic results exist only in a few special cases [kitanine1999form, kitanine2000correlation, boos2007factorization, caux2007one, trippe2010short, pozsgay2011local, negro2013one, mestyan2014short, piroli2015exact, piroli2016multiparticle, pozsgay2017excited, bastianello2018exact, bastianello2018sinh].
@@
The picture is very different for JW free fermions, where a number of simplifications occur. First, the JW transformation maps local bilinear fermionic operators into local spin operators. Second, it is often easy to characterize the family of spin states that map to the fermionic Gaussian states [^1] We recall that Gaussian states are defined as those satisfying the Wick theorem [bravyi2004lagrangian, surace2022fermionic].. These facts have consequences for the complexity of the JW solvable models. For instance, the dynamics of spin correlation functions can be computed either analytically or numerically efficiently in many cases of interest [terhal2002classical, vanDenNest2011simulating, brod2016efficient], and especially in the case of quantum quenches [calabrese2006time, fagotti2008evolution, calabrese2011quantum, essler2016quench]. In addition, the eigenstate wave functions are simple compared to generic Hamiltonians. In particular, all eigenstates can be prepared efficiently by a quantum computer, as they can be obtained as the output of a quantum circuit whose depth scales only linearly in the system size [verstraete2009quantum, ferris2014fourier, kivlichan2018quantum, jiang2018quantum, hackl2018circuit]. It is natural to ask whether the same conclusions can be drawn for FFD models. Can the system dynamics be simulated on a classical computer? Can all eigenstates be prepared efficiently on a quantum computer? [lutz2025adiabatic] These questions are non-trivial, due to the complicated structure of the dictionary between spin and fermion degrees of freedom.
@@
In this work, we take a step towards answering these questions, by characterizing the Hilbert space associated to FFD Hamiltonians. For concreteness, we will focus on the original model introduced by Paul Fendley [fendley2019free], although similar constructions hold for more general FFD Hamiltonians. We show that the corresponding Hilbert space admits the exact factorization $\mathcal{H}=\mathcal{H}_{F}\otimes\mathcal{H}_{D}$, where $\mathcal{H}_{F}$ hosts the fermionic operators, while $\mathcal{H}_{D}$ accounts for the exponential degeneracy of the energy eigenspaces. By constructing a family of spin operators generating the operator algebra supported on $\mathcal{H}_{D}$, we further show that $\mathcal{H}_{D}=\mathcal{H}_{F^{\prime}}\otimes\mathcal{H}_{\widetilde{D}}$, where $\mathcal{H}_{F^{\prime}}$ hosts ancillary free fermions in disguise, while $\mathcal{H}_{\widetilde{D}}$ is generated by the common eigenstates of an extensive set of commuting Pauli strings. Our construction allows us to fully resolve the exponential degeneracy of all Hamiltonian eigenspaces and is expected to have implications for the computation of spin correlation functions, both in and out of equilibrium.
@@
This work is organized as follows. We begin in Sec. II by recalling the model introduced by Paul Fendley (Sec. II.1) and its exact solution in terms of free fermions (Sec. II.2). Our results are contained in Secs. II.3, IV, and V, and summarized in Sec III. In Sec. II.3 we show that the Hilbert space factorizes as $\mathcal{H}=\mathcal{H}_{F}\otimes\mathcal{H}_{D}$, where $\mathcal{H}_{F}$ supports the fermionic operators. The structure of $\mathcal{H}_{D}$ is further resolved in the later sections, where we show that $\mathcal{H}_{D}=\mathcal{H}_{F^{\prime}}\otimes\mathcal{H}_{\widetilde{D}}$ and characterize the two factors individually. In particular, in Sec. IV we show that $\mathcal{H}_{\widetilde{D}}$ is a “representation-dependent” subspace, generated by the common eigenstates of an extensive set of commuting Pauli strings, while in Sec. V we show that the space of operators supported on $\mathcal{H}_{F^{\prime}}$ is generated by an operator algebra of ancillary free fermion in disguise. Finally, Sec. VI contains our conclusions and an outlook for future work.
II The Model and its Hilbert-space structure
@@
II.1 The Hamiltonian
We first introduce the FFD Hamiltonian presented in Ref. [fendley2019free]. The construction starts with an abstract operator algebra defined in terms of a set of generators $\{h_{m}\}_{m=1}^{M}$. They satisfy the commutation relations
@@ II.1 The Hamiltonian
Different representations of this algebra give rise to different models. In this work, we focus on the original representation considered in Ref. [fendley2019free], which is associated with a one-dimensional chain of $M$ spins. The Hilbert space of the model is $\mathcal{H}=(\mathbb{C}^{2})^{\otimes M}$, while
@@
where $b_{m}\in\mathbb{R}$ are arbitrary real numbers parametrizing the model (note that the Hamiltonian (2) corresponds to open boundary conditions). In the next subsection, we review the exact solution of the Hamiltonian (2). We will focus on the aspects that are directly relevant to our work, and refer to Ref. [fendley2019free] for more detail.
II.2 Exact solution
The diagonalization of the Hamiltonian (2) is achieved by constructing suitable fermionic creation and annihilation operators. A key object is the transfer matrix $T_{M}(u)$, an operator acting on $\mathcal{H}$ that depends on a complex parameter $u$, called the spectral parameter. The transfer matrix can be constructed as the generating function of the conserved quantities of the model [fendley2019free], in a fashion which is reminiscent of Bethe-ansatz solvable models [korepin1997quantum].
@@ II.2 Exact solution
As shown in [fendley2019free], the creation/annihilation operators associated with the energy $\epsilon_{k}$ then take the form
$\Psi_{k}\propto T_{M}(\epsilon_{k})\penalty 10000\ \chi\penalty 10000\ T_{M}(-\epsilon_{k})\,.$ (7)
@@
II.3 The Hilbert-space structure
In this section, we use the fermionic operators (7) to show that the FFD Hilbert space can be decomposed as $\mathcal{H}=\mathcal{H}_{F}\otimes\mathcal{H}_{D}$. Here, $\mathcal{H}_{F}$ is a space of dimension $2^{S}$ that the fermionic operators are supported on, while $\mathcal{H}_{D}$ accounts for the exponentially large degeneracies observed in the spectrum of $H$. This simple observation is our first result and the basis for the subsequent analysis presented in the rest of this work, cf. Sec. III
@@ II.3 The Hilbert-space structure
[…]
@@
[…]
III Overview of our results
@@
  • We will show that ${\mathcal{H}}_{F^{\prime}}$ supports the action of all operators written in terms of the generators $h_{m}$ that commute with the fermion creation/annihilation operators. We will denote the algebra of such operators by $\mathcal{A}_{F^{\prime}}$ and we will refer to it as the ancillary fermionic algebra (to be distinguished by the fermionic algebra $\mathcal{A}_{F}$ supported on $\mathcal{H}_{F})$. Importantly, the definition of $\mathcal{A}_{F^{\prime}}$ is independent of the specific representation chosen, because it can be stated entirely in terms of the generators $h_{m}$ and their algebraic relations. However, we will show that $\mathcal{A}_{F^{\prime}}$ does depend on the Hamiltonian couplings $\{b_{m}\}$. Going further, we will show in Sec. V, that $\mathcal{A}_{F^{\prime}}$ is a fermionic algebra, generated by a new family of ancillary fermion operators
@@
The dimensions of the subspaces $\mathcal{H}_{F}$, $\mathcal{H}_{F^{\prime}}$, and $\mathcal{H}_{\widetilde{D}}$ depend, respectively, on the integers $S$, $S^{\prime}$, $N_{P}$ and $N_{C}$ introduced above. In Table 1, we summarize their dependence on the system size $M$. In all cases (except for the case $M=6k+3$, where an additional zero mode has to be included to the construction, cf. the end of Sec. V), we can check that the product of the dimension of the subspaces $\mathcal{H}_{F}$, $\mathcal{H}_{F^{\prime}}$, and $\mathcal{H}_{\widetilde{D}}$ is equal to the dimension of the full Hilbert space, ${\rm dim}(\mathcal{H})=2^{N}$.
@@
$M$ $S=\lfloor\frac{M+2}{3}\rfloor$ $N_{C}$ $N_{P}$ $\penalty 10000\ S^{\prime}=\lfloor\frac{M+2}{6}\rfloor$
@@
It is important to compare our results with some constructions found in Ref. [fendley2019free]. There, a family of operators $E^{(n)},O^{(n)}$ is presented, commuting with the Hamiltonian and generating an extended supersymmetry (SUSY) algebra. While this symmetry also accounts for the degeneracies of the spectrum, it has a few drawbacks. First, it is not clear what the underlying algebra is. Second, the generators $E^{(n)},O^{(n)}$ mix with the fermions: namely, they entangle the spaces $\mathcal{H}_{F}$ and $\mathcal{H}_{D}$, which we believe makes it difficult to understand the corresponding algebra  [^2] In our construction, we will also encounter operators $E^{(n)},O^{(n)}$. Those are defined much similarly to Fendley’s, but with the important difference that their relation with the fermionic operators is clearly understood..
IV Representation-dependent degeneracies: the subspace $\mathcal{H}_{\widetilde{D}}$
As pointed out in Ref. [fendley2019free], the representation (1) allows for a family of operators $\{\widetilde{h}_{m}\}$ that commute simultaneously with all the $h_{m}$, and therefore generate a symmetry of the model which does not depend on the Hamiltonian couplings $\{b_{m}\}$. More generally, such symmetry extends to any type of quantum dynamics defined in terms of the operators $\{h_{m}\}$, such as the circuits considered in Refs. [fukai2025quantum, szasz2025construction]. The operators $\widetilde{h}_{m}$ can be chosen as [fendley2019free]
@@ IV Representation-dependent degeneracies: the subspace $\mathcal{H}_{\widetilde{D}}$
each pair $(\tilde{X}_{n},\tilde{Y}_{n})$ accounting for a two-fold degeneracy of the Hamiltonian spectrum. For some values of $M$ modulo $6$, these two families are completed by either one or two additional central operators, commuting with all $\tilde{X}_{m}$ and $\tilde{Y}_{m}$. Of these central operators, some belong exclusively to the algebra of the $\widetilde{h}_{m}$, while others can also be formulated as elements of the algebra of the $h_{m}$ (see paragraphs IV.5 and IV.6): for reasons that will be made clear below, the index $N_{C}$ in Table 1 counts only the former.
IV.1 $M=6k$